(8) Find the least-squares solution for Az = 5 where 5 and 4 1 A= 2 -1 -2 0 -3 2 -5
Answer #1

anteet sa 6.45 -6 8) Least squares Soltion A2 = 6. 3. -3 b Given A - 2 -3 2-5 . Since Αχ -και AT on both sides) Using ATAR =6 - 6 x asked Since solution is -6 - 6x +42y = -6 62-by. -60 ΑΣ - 4x +424.-6 62-64..6. 031 3 by -12 1-6 62-6 (7/3) = -6 4:62